OpenStudy (anonymous):

write the expression using only natural logarithms : log base2 x

OpenStudy (anonymous):

Well, \[\log_ab = \frac{\ln{b}}{\ln{a}} \] so that would just be \[\log_2x = \frac{\ln x}{\ln 2} \]

OpenStudy (anonymous):

Wow I feel dumb. Thanks :D